You measure an electric field of 2.52×106 N/C at a distance of 1.196 m from a point charge. What is the magnitude of the charge?
1
Expert's answer
2020-02-27T09:48:23-0500

Solution. The electric field of a point charge can be found behind the formula


"E=\frac{kQ}{r^2}"

where r=1.196 m is a distance; Q is charge; k=9×10^9 N m2C-2 is Coulomb's constant.

Therefore


"Q=\frac{Er^2}{k}=\frac{2.52\times 10^6\times 1.196^2}{9\times10^9}\approx 4\times10^{-4}C"

Answer.

"Q=4\times10^{-4}C"